OpenStudy (anonymous):

find the inverse of each matrix. [2 -5] 9] [5

OpenStudy (zzr0ck3r):

\[\left[\begin{matrix}2 & -5 \\ 5 & 9\end{matrix}\right]?\]

OpenStudy (zzr0ck3r):

\[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]^{-1}= \frac{1}{ad-cb}\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (zzr0ck3r):

@fabo ?